2. Of the 38 plays attributed to Shakespeare, 18 are comedies, 10 are tragedies, and 10 are histories. One play is randomly selected from Shapespeare’s 38 plays. Find the odds.

a. In favor of selecting a comedy.


b. In favor of a selecting a tragedy.


c. Against selecting a history.


d. Against selecting a comedy.


e. In favor of selecting a comedy or a tragedy.



f. In favor of selecting a tragedy or a history.



g. Against selecting a tragedy or a history.



h. Against selecting a comedy or a history.
1
Expert's answer
2020-05-06T19:02:45-0400

"(a)\;\frac{18}{38-18}=\frac{18}{20}=0.9\\(b)\;\frac{10}{38-10}=\frac{10}{28}=\frac5{14}=0.357\\(c)\;\frac{38-10}{10}=\frac{28}{10}=\frac{14}5=2.8\\(d)\;\frac{38-18}{18}=\frac{20}{18}=\frac{10}9=1.111\\(e)\;\frac{18+10}{38-10-18}=\frac{28}{10}=\frac{14}5=2.8\\(f)\;\frac{10+10}{38-10-10}=\frac{20}{18}=\frac{10}9=1.1\\(g)\frac{38-10-10}{10+10}=\frac9{10}=0.9\\(h)\frac{38-10-18}{10\;+\;18}=\frac5{14}=0.357"